In an integrated circuit, the current density in a

2.2-μm-thick × 80-μm-wide gold film is 7.9×105 A/m2 .

How much charge flows through the film in 15 min?
Answer #1

Current density =J= 7.9*10^5 A/m^2

Area = A=2.2*80 = 176*10^-12 m^2

Time = t = 900 sec

Charge = Q = I*t = J*A*t = 0.125136 C